Circular No. 6166

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995M IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova (magnitude R ∼ 18.5) on a plate taken by C. P. Cass with the U.K. Schmidt Telescope on Apr. 22.41 UT. SN 1995M is at α = 9^{h}38^{m}41^{s}.78, δ = −12^{o}20′07′′.9 (equinox 2000.0, uncertainty in both coordinates 0′′.3), which is 13′′ due west of the center of the edge-on host galaxy, near the tip of the disk as recorded on the deep survey plate. No star appears at this position on earlier R or J exposures or on I plates taken as late as 1994 Mar. 18. A nearby star of mag ∼ 19 has position end figures 41^{s}.84, 19′30′′.9. P. Garnavich, A. Riess, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that P. Zhao obtained a CCD image on Apr. 24.3 UT with the 1.2-m Mt. Hopkins telescope showing SN 1995M as a star of mag V = 18.2 ± 0.1 at the western tip of an edge-on galaxy; it does not appear on the Palomar Sky Survey prints down to a magnitude limit of ∼ 20.

B. Schmidt, Mount Stromlo and Siding Spring Observatories; and R. Stathakis, Anglo-Australian Observatory, obtained a spectrogram (range 360–740 nm) on Apr. 24.5 at the Anglo-Australian Telescope that shows this object to be a type-Ia supernova ∼ 15 days after maximum. The redshift of the host galaxy, measured from narrow emission lines at the position of SN 1995M, is z = 0.052.

SUPERNOVA 1995D IN NGC 2962 H. Mikuž, Ljubljana, reports additional CCD photometry (cf. IAUC

6149) of SN 1995D: Mar. 11.94 UT, V = 14.23 ± 0.03, B − V = +0.79 ± 0.05; Apr. 3.78, 15.39 ± 0.03, +1.25 ± 0.05; 8.81, 15.59 ± 0.03, +1.12 ± 0.06; 17.81, 15.81 ± 0.03, +1.05 ± 0.06.

WORLD WIDE WEB The Central Bureau has recently inaugurated a homepage on the World

Wide Web giving online information about the services offered by the Bureau. Computer-service subscribers may access the Central Bureau computers via telnet through this homepage. The Uniform Resource Locator

for this homepage is http://cfa-www.harvard.edu/cfa/ps/cbat.html.

1995 April 25 Daniel W. E. Green